OpenStudy (anonymous):

7. What is the value of 7P3? with permutations

OpenStudy (anonymous):

\[7\times 6\times 5\]

OpenStudy (anonymous):

210

OpenStudy (anonymous):

840?

myininaya (myininaya):

210 is right

myininaya (myininaya):

\[n P r=\frac{n!}{(n-r)!}\]

myininaya (myininaya):

that is the formula

myininaya (myininaya):

\[7 P 3=\frac{7!}{(7-3)!}=\frac{7!}{4!}=\frac{7 \cdot 6 \cdot 5 \cdot 4!}{4!}=7 \cdot 6 \cdot 5 \]